What capacitance is required to store an energy of 7.9kW·h at a potential difference of 1100 V?
Help Please!!
Thanks:)
Answer #1
Energy, E = 7.9 kW-h = 7.9*1000*3600 W.s (or J) =2.844x107 J
Potential difference, V = 1100 V

Now, Energy stored in capacitor, E = (1/2)CV2,
which gives, 2.844x107 = (1/2)*C*(1100)2, andwe get, C = 47.01 Farad
Answer #2
Energy(W)=7.9*1000*3600 J =2.84x107 J
W=1/2CV2
C=2W/V2
   =2*2.84x107/11002
   
=47.01 F